Among all pairs of numbers with a sum of 139, find the pair whose product is maximum. Write your answers as fractions reduced to lowest terms.
Answer #1

2 Let the numbers whose sum is 139 ane x and y. :: x+y = 139 => y = 139-2 .. Producer = rey = x(1.39 ~~) - = 139x-x2 = -32713